Telescope and AGw Commissioning at LFBG with IRTC - 07-08 April 2009 MST

Observers: DThompson, MPedani (until 2:00am), ARakich (from midnight) (in Tucson), JHill (@LBT until 8:00pm)
Telescope Operator: D. Gonzalez Huerta
Telescope Support: J. Little

Summary

Lost night to bad weather. 100% cloudy first half, 22+m/s wind second half. Called it a night at 2:30AM to prepare for possible laser-tracker and calibration unit tests of AGW on telescope tomorrow. Weather forecast for tomorrow night is now 40% chance of snow.

Details

18:32 Weather on mountain does not look good for tonight.

Added +35 arcsec RX to IRTC M1 collimation model to correct for some initial coma. We need to initialize on PSF to load the change.

Modified IRTC pointing model to include -70 arcsec offset to IE due to collimation model RX change. The conf file was changed to make this new pointing model (dated 20090408) the default. Need to reload pointing model on PCS to implement the change.

JMH: Later in the day, I applied Andrew's -3.5 arcsec tip to both the IRTCF and LUCIFER secondary collimation models SXSMLUCIFERCollimation.dat and SXSMIRTCFCollimation.dat on 08-APR-2009. This also moved IE by +4 arcsec.

02:30 Tair=2.3C, RH=54%, Wind=22m/s gusts

-- JohnHill - 07 Apr 2009
Topic revision: r5 - 09 Apr 2009, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback